OpenStudy (anonymous):

Find the measure in radians of the central angle of a circle with radius of 4 inches and arc length 8 inches. Options are 32, 0.5, 2, or 2pi.

OpenStudy (tkhunny):

What is the arc length of the entire circle?

OpenStudy (anonymous):

This is all the information given.

OpenStudy (tkhunny):

It's a circle of radius 4. I ask again, what is the arc length of the entire circle? It doesn't have to be in the problem statement. You should be able to calculate it with the given information.

OpenStudy (anonymous):

Is the arc length of the entire circle 16? (2*8), or 32? (8*4)?

OpenStudy (tkhunny):

\(2\pi r = 8\pi\;inches\) \(\dfrac{8\;inches}{8\pi\;inches} = \dfrac{Internal\;Angle}{2\pi}\)